OpenStudy (anonymous):

he number line below represents which combined inequality? number line with a closed circle on -4, shading to the left and a closed circle on 3, shading to the right x greater than or equal to -4 and x less than or greater to 3 x greater than or equal to -4 or x less than or greater to 3 x less than or greater to -4 and x greater than or equal to 3 x less than or greater to -4 or x greater than or equal to 3

OpenStudy (anonymous):

@AndrewTheCookie

OpenStudy (anonymous):

Someone please help

OpenStudy (blacksteel):

When you say "shading", do that mean that there is a thick line going in that direction?

OpenStudy (anonymous):

I do not know @Blacksteel it just was like that

OpenStudy (anonymous):

that is what my teacher is asking for

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@KamiBug

OpenStudy (kamibug):

A closed circle means it is also equal to. If it is shaded to the left it means less than. An open circle means not equal to. If it is shaded to the right it means greater than. ...So what do you think? :)

OpenStudy (anonymous):

answer is b? srry not very good at this

OpenStudy (kamibug):

It would be choice B if the line from 3 was shaded to the left. But, it is shaded to the right. So it should be options C or D, if they say "x is less than or equal to -4 ... x is greater than or equal to 3" I'm not sure what it means by the "and" "or" Sorry. :/

OpenStudy (anonymous):

or is meaning the less than or greeter to signs it has the line underneath the inequality sign @KamiBug

OpenStudy (anonymous):

or means there is an option and means there is not

OpenStudy (anonymous):

@Blacksteel may you also help me

OpenStudy (kamibug):

The line underneath the sign means "or equal to" I'm thinking the "and"/"or" has something to do with the space between the two shaded lines on the number line, but not so sure...

OpenStudy (anonymous):

ok hold on please

OpenStudy (kamibug):

Okay :)

OpenStudy (anonymous):

i do not know to be honest

OpenStudy (anonymous):

may you just help me with another

OpenStudy (kamibug):

I'd go with the "or" :P Try D... :)

OpenStudy (anonymous):

thank you may you help me with a few more?

OpenStudy (kamibug):

Sure :)

OpenStudy (anonymous):

Henry's rocket is designed to launch to a height of 100 feet in the air. The measurements so far have the launch height (h) varying no more than 4 feet. Which inequality below represents this scenario? |h - 4| less than or greater to 100 |h + 4| greater than or equal to 100 |h - 100| less than or greater to 4 |h + 100| greater than or equal to 4

OpenStudy (anonymous):

I am thinking this one is c

OpenStudy (kamibug):

I agree. ^.^

OpenStudy (anonymous):

Which of the following is not equivalent to the formula d = rt? d over r= t r = dt d over t= r d over r t= 1

OpenStudy (kamibug):

r = dt is incorrect. It should be r = d/t

OpenStudy (anonymous):

why do you feel that

OpenStudy (kamibug):

Well, to solve d = rt for r, you need to divide both sides by t, not multiply...

OpenStudy (anonymous):

well they are asking what is not correct not what is correct

OpenStudy (anonymous):

After a discount of 10%, the price of a pair of jeans, including tax, was $69.62. The following equation can be used to find x, the price of the pair of jeans without the discount: x - 0.1x + 0.07(x - 0.1x) = 69.62 What is the price of the pair of jeans without the discount? $70.58 $72.29 $74.86 $80.02

OpenStudy (kamibug):

Yeah, B is the answer because it is incorrect. That's what I mean. :P

OpenStudy (anonymous):

ok thanks @KamiBug

OpenStudy (anonymous):

please help with the question I just asked please

OpenStudy (anonymous):

@KamiBug

OpenStudy (anonymous):

@Blacksteel may you help me with the one I just asked Kami please

OpenStudy (blacksteel):

Sure. This problem requires you to distribute a coefficient, combine like terms, and then isolate the variable. Let's start with our given equation. \[x - 0.1x + 0.07(x - 0.1x) = 69.62\]Now we're going to distribute the coefficient of the terms in parentheses to the individual terms and simplify. \[x - 0.1x + 0.07x - 0.07*0.1x) = 69.62\]\[x - 0.1x + 0.07x - 0.007x) = 69.62\]Now we want to combine like terms; since all of the terms are multiples of x, we can add them together. \[0.963x = 69.62\]Finally, we want to isolate x by dividing both sides of the equation by the coefficient of x. \[0.963x/0.963 = 69.62/0.963\]\[x = 72.29491...\]To get our answer, we simply round: x = $72.29

OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality: 2x + 5 > 7 and 5x less than or greater to 25 A number line with an open circle on 1, shading to the left, and a closed circle on 5, shading to the right A number line with an open circle on 1, a closed circle on 5, and shading in between A number line with a closed circle on 1, shading to the left, and an open circle on 5, shading to the right A number line with a closed circle on 1, an open circle on 5, and shading in between

OpenStudy (loser66):

5x less than or greater to 25 means x less than or greater to 5. That is x is everything but 5

OpenStudy (loser66):

2x+5 >7 2x >2 x> 1  so, combine them, you have ??? You should try once. I can check. Please, don't guess.

OpenStudy (anonymous):

The answer is c. @Loser66

OpenStudy (loser66):

nope, the last one

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

Sean, a freelance editor, charges the rates shown in the table below to edit manuscripts. The cost per page increases as the quality of editing improves. Sean also gives a 5% discount if the entire amount is paid up front. A 2-column table with 'Type of Editing' and 'Cost per page' as the column headings. Express Proofreading costs $2.00, Basic Proofreading - $2.95, Extended Proofreading - $3.95, and Deep Editing - $11.00 Michelle has an 85-page manuscript. The equation below shows the relationship between the total cost of editing 85 pages, T, and the cost per page, c, if she gets the 5% discount: 85c - 0.05(85c) = T Using the equation, what is the best quality of editing that Michelle can get done for a maximum of $161.50? Express proofreading Basic proofreading Extended proofreading Deep editing

OpenStudy (anonymous):

@texaschic101